You have just been hired as a new management trainee by Earrings Unlimited, a distributor of earrings to various retail outle749,600 152,768 29,000 1,110,000 $2,131,368 $691, 200 March sales) Inventory Prepaid insurance Property and equipment (net) TThe company has an agreement with a bank that allows the company to borrow in increments of $1.000 at the beginning of each mThe company has an agreement with a bank that allows the company to borrow in increments of $1.000 at the beginning of each mThe company has an agreement with a bank that allows the company to borrow in increments of $1.000 at the beginning of each m
Answer #1

Calculation April May June Earrings Unlimited Sales Budget Req la Sales in Units Sales price per unit Total Sales April May 6April 1,364,000 May 2,064,000 June May June 1,064,000 April Quarter 4,492,000 =68,200*5.60 1,364,000*4% =103,200*5.60 2,064,0